All Systems Operational
Coventry - Continuity House (205/DC3) Operational
90 days ago
99.99 % uptime
Today
Coventry (DH1) Operational
90 days ago
100.0 % uptime
Today
Coventry (DH2) Operational
90 days ago
100.0 % uptime
Today
Telehouse East (THE) Operational
90 days ago
100.0 % uptime
Today
Telehouse North (THN) Operational
90 days ago
99.99 % uptime
Today
Telehouse West (THW) Operational
90 days ago
100.0 % uptime
Today
Pulsant 2 (MHD) Operational
90 days ago
100.0 % uptime
Today
Equinix MA1 (MCR) Operational
90 days ago
100.0 % uptime
Today
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Major outage
Partial outage
No downtime recorded on this day.
No data exists for this day.
had a major outage.
had a partial outage.
Past Incidents
May 18, 2024

No incidents reported today.

May 17, 2024
Resolved - This incident has been resolved.
May 17, 14:59 BST
Update - This wavelength service has been restored and we have bought this link back into use again across our network.

UK Servers customers should not have noticed any impact to their services during the outage of this circuit and our Coventry site is no longer deemed at risk.

May 17, 01:30 BST
Update - The cable pull has now been completed and splicers have begun their work.

Please note, this Neos Networks NOC informs us this is a 216 fibre cable and repairs are likely to continue throughout the night.

May 16, 23:00 BST
Update - The fault has been caused by National Grid installing an new pylon, work has commenced to replace the damaged duct, once this has been completed replacement fibre will need to be pulled and re-spliced to restore service.
May 16, 18:11 BST
Update - Engineers have completed OTDR testing to measure where the fault on the fibre is, they are now commencing testing to pin point the exact location on the break.

We have no estimated time of restoration as of yet.

May 16, 15:23 BST
Identified - The latest update from the provider of this circuit is as follows:

We are currently investigating a major incident affecting multiple services due to a suspected fibre break between Evesham and Bradley Stoke.

Fibre engineers have been dispatched to pin point the exact location of the break.

Updates to follow once engineers have arrived on site and started their investigation.

May 16, 13:14 BST
Investigating - At approximately 11:58 a wavelength circuits between our Continuity House Coventry Datacentre and THW (Telehouse West) that provides connectivity to this site went offline and remains down.

Due to the redundant nature of our network this has not impacted customers however pleased be advised that whilst this circuit remains offline the network at our Coventry sites is deemed "at risk"

May 16, 12:15 BST
Resolved - We have recieved the all clear from the Colt NOC and this wavelength service has been restored. Our Coventry datacentre is no longer deemed at risk.
May 17, 00:30 BST
Monitoring - The field technician replaced a faulty card on site, and all alarms have been clear since 23:28, we are continuing to monitor this circuit.
May 16, 23:35 BST
Identified - The field technician is on site and will be starting the emergency repair work. We will keep you posted as per progress on restoration work.
May 16, 22:28 BST
Investigating - We have been informed by the Colt NOC that they have detected faulty hardware, causing unprotected customer impairment and leaving services at risk of failure. As such, they have implemented an urgent maintenance window to fix optical cables and prevent future service interruptions. A 2-hour maintenance window has been scheduled for this activity starting at 22:00 on 16th May 2024.

Due to the redundant nature of our network this should not impact customers however pleased be advised that the network will be at a high risk level during this work particularly as other wavelength services connecting to our Coventry site are also currently offline.

May 16, 21:27 BST
May 16, 2024
May 15, 2024

No incidents reported.

May 14, 2024

No incidents reported.

May 13, 2024

No incidents reported.

May 12, 2024

No incidents reported.

May 11, 2024

No incidents reported.

May 10, 2024
Completed - The scheduled maintenance has been completed.
May 10, 06:00 BST
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
May 9, 22:00 BST
Scheduled - During the above period one of the suppliers of wavelength services providing connectivity from our London PoPs to our Coventry datacentres will be carrying out essential maintenance work.

This work involves migrating traffic away from the CPL equipment at site.

Due to the redundant nature of the UK Servers network this maintenance is not expected to be service affecting for our customers as traffic will automatically re-route on alternative paths however due to the nature of the work the network to our Coventry site should be deemed at risk during this period.

Whilst a lengthy maintenance window is scheduled to cover all eventualities the service provider does not expect the circuits to be offline for the duration of the maintenance window and UK Servers customer's services should not be affected.

Apr 25, 14:22 BST
May 9, 2024
May 8, 2024

No incidents reported.

May 7, 2024

No incidents reported.

May 6, 2024

No incidents reported.

May 5, 2024

No incidents reported.

May 4, 2024

No incidents reported.